[figure: slide left; jump right; slide right; jump left; jump left; progn; slide left; jump right; jump right; jump right; slide right; jump left; jump left; progn; jump left; jump left; progn; slide left; jump right; jump right; jump right; jump right; progn; jump right; progn; slide left; jump left; jump left; jump left; jump left; slide right; jump right; jump right; jump right; progn; slide left; jump left; jump left; progn; slide right; progn; jump right; progn; progn; slide left; slide left]

1